To keep a vehicle moving at a constant speed "V" require power "P" from the engine. The force provided by the engine is?


1
Expert's answer
2022-01-07T09:29:43-0500

The power by definition


"P=\rm\frac{work\: done}{time}"

"P=\frac{W}{t}=\frac{Fs}{t}=Fv"

Hence, the force

"F=\frac{P}{v}"